TMS Meeting Minutes, September 30, 2021

Attendees: Andrew, Bo Xin, Breann, Heejoo, John, Olga, Trenton, Yang (Zoom Meeting)

Updates:
  • Leroy did a great job re-aligning Ch. 10 and Ch. 21, as indicated by the channel strength at 51 deg elevation
    • Ch. 10 was the only channel in use that is below 10% (around 4%)
    • We can still increase the gain on Ch. 10 to include this
    • At 30 degree elevation, there was some dropoff in the signal strength, but the gains could be adjusted to compensate
  • Leroy also improved the collimator mounting to reduce rocking
  • Leroy also checked the pairing between the RR and the collimators and was unable to find any disparities between our dataset from the last visit
  • Heejoo and John did some more testing to determine that the system can be further optimized to work at 51 deg and 30 deg
  • One thing that was discovered today is that we have a repeatable pattern, where the TMS throws NaN's on the Sx side, followed by NaN's on the other side
    • This appears to be most likely when tracking a single OB while changing filters with 1 sec exposures
    • So long as TMS does no harm when it fails a measurement during this sort of a drop
  • Andrew wants to call attention to the pointing issues that John brought up a few days ago
    • It appears that there are problems with the LBT pointing control loops that had not yet been discovered
    • TMS may provide insight into this issue
  • See John's detailed summaries on Google Drive (saved email correspondence) for the details on the recent findings
    • John notes that there appears to be a leak in the range balancing code
    • There's an offset being applied; the combination of multiple range balancing contributions seems to cause drift away from collimation
    • Currently, there is a gradual ramp-up in the offset, as a continuous error that adds up over time
  • 20210914 TMS Right Plot
    • There are steps in pointing correction which don't exactly correspond with the time that slewing to the new OB occurs
  • Now that we have the stabilization of the collimators addressed, Andrew thinks it is time to revisit the start-of-the-night set
    • Use as a starting reference the reference from the last night, before FPIA
    • Able to measure the structure change and measure a focus change using a previous reference
    • The thermal focus is currently in the code, but the thermal focus references itself to the hour of the last set_ref run (so this won't work well when using a reference from the past)
    • John simply needs to make the code smarter to not apply the thermal focus correction if the reference was taken from the day prior
  • Heejoo and Trenton were originally planning to head up the mountain next week, but there is no need for us to go up now with the resolution of the hardware issues
  • Possibility of a TMS opportunity tonight, following LUCI checkout operations
    • Weather might present an issue, and LUCI operations won't be happening if seeing is worse than 1.2 arcsec
    • Andrew would like to run science like observations
    • John planning to take one standard star observation to prove we have overcome our recent issues, then move to an hour on an OB without running FPIA to prove consistent functioning of TMS
    • Useful to make note of "non-photometric"
  • Trenton still making progress on the primary mirror modeling
    • John notes that it is very important to consider the gravity sag induced by poorly chosen constraints
    • This parabolic sag departure can completely overshadow other effects if not accounted for correctly with realistic actuator forces (too complex)
    • For simplicity, it may be easiest to just focus on the thermal effects without considering the gravitational effects
-- TrentonBrendel - 30 Sep 2021
Topic revision: r1 - 30 Sep 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback